OpenStudy (anonymous):

A motorist driving at 25 meters/second decelerates to 15 meters/second. The combined mass of the vehicle and the driver is 1.5 × 10³ kilograms. What is the amount of mechanical work done by the vehicle?

OpenStudy (anonymous):

Work done = change in energy = 0.5*m*v1*v1 - 0.5*m*v2*v2 \[\frac{ 1 }{ 2 } m \left( v_i^2 - v_2^2 \right) = \frac{1}{2} 1.5 \times 10^3 \left(25^2 - 15^2 \right) = 300 \times 10^3 N\]

OpenStudy (anonymous):

i get it now, but how would i work this out? An object having a mass of 2.0 kilograms falls from a height of 15 meters. What is its kinetic energy when it hits the ground?

OpenStudy (anonymous):

Gain in kinetic energy = loss in potential energy = m*g*change in height \[mg \Delta h = 2 * 9.8 * \left( 15 - 0 \right) = 294\]

OpenStudy (anonymous):

thanks, i just passed my test haha

mathslover (mathslover):

@lycan1995 though I had made this aware to admins/mods but I am still requesting you : This site only helps those people who are keen to help themselves first.. I actually mean to say that you must have tried that first... Oh k so leave this now,, the main point is: This site does not help the users fort the tests... they should attempt them themselves... I hope you got my point though if you need more information consider: http://openstudy.com/code-of-conduct Also since you are a new user : Welcome to this world of education \(\huge{\color{blue}{\cal{OpEnStUdY}}}\)

myininaya (myininaya):

@lycan1995 I have to agree with @mathslover . Please don't use OpenStudy to cheat on your tests in the future.